OpenStudy (anonymous):

z^-7/z^-4 show step how to solve

OpenStudy (anonymous):

So in this problem you would need to subtract the exponents since they have the same base which is z.

OpenStudy (anonymous):

ok i got 2^3

OpenStudy (anonymous):

@200205650

OpenStudy (anonymous):

Where did you get a 2? The base is a z.

OpenStudy (anonymous):

oh my bad i ment z^3

OpenStudy (anonymous):

\[\frac{ z ^{-7} }{ z^{-4} }\]Take the exponents. -7 & -4 Now subtract them. -7-(-4) = -3 -3 is now your exponent. \[z^{-3}\]But negatives can't be with exponents so in order to get rid of it, you would have to flip the whole "z^-3". The final answer should look like this: \[\frac{ 1 }{ z^{3} }\]

OpenStudy (anonymous):

You were close.

OpenStudy (anonymous):

Did you understand my explanation?